A neon light is turned on, emitting photons that are 700 nm long and travels at C. What is the EMR’s: a) frequency

b) period and c) what color of visible light is it?
1
Expert's answer
2020-10-30T06:41:37-0400

"E=hf"

"E=hf=" "hc\over \lambda"

Where "E=700nm"

"h=4.14\times 10^{-15}eV.s."

"hc=1240eV.nm"

a) Frequency(f)

"700 nm=" "4.14\times 10^{-15}eV.s. \times (f)"

"f=" "700nm\over 4.14\times 10^{-15}eV.s."

"=1.690\times 10^{17} nm^{-1}eV.s^{-1}"

b) From the equation"E=" "hc\over \lambda"

Where "hc=1240eV.nm"

"\therefore \lambda=" "1240eV. nm\over700nm" "=1.771eV"

The frequency is therefore "1.771eV"

c) From the "EM" spectrum, the color of the neon light is Ultraviolet.